# Multi-Head RAG: Solving Multi-Aspect Problems with LLMs

### v2

@@ Abstract
Retrieval Augmented Generation (RAG) enhances the abilities of Large Language Models (LLMs) by enabling the retrieval of documents into the LLM context to provide more accurate and relevant responses. Existing RAG solutions do not focus on queries that may require fetching multiple documents with substantially different contents. Such queries occur frequently, but are challenging because the embeddings of these documents may be distant in the embedding space, making it hard to retrieve them all. This paper introduces Multi-Head RAG (MRAG), a novel scheme designed to address this gap with a simple yet powerful idea: leveraging activations of Transformer’s multi-head attention layer, instead of the decoder layer, as keys for fetching multi-aspect documents. The driving motivation is that different attention heads can learn to capture different data aspects. Harnessing the corresponding activations results in embeddings that represent various facets of data items and queries, improving the retrieval accuracy for complex queries. We provide an evaluation methodology and metrics, multi-aspect datasets that we release online, and real-world use cases to demonstrate MRAG’s effectiveness, showing improvements of up to 20% in relevance over standard RAG baselines. MRAG can be seamlessly integrated with existing RAG frameworks and benchmarking tools like RAGAS as well as different classes of data stores.
@@ 1 Introduction
Large Language Models (LLMs) transformed many machine learning tasks using in-context learning abilities. They achieved such accuracy by leveraging an increasing number of parameters, which in recent models have grown to hundreds of billions, making LLM training expensive in terms of both time and resources. It also comes with the danger of leaking confidential data into model weights (Yan et al., 2024; Shi et al., 2024; Patil et al., 2024). Additionally, continuous training through fine-tuning is necessary to keep LLMs up-to-date. Even using the newest data, LLMs display an ongoing problem of hallucinations (Zhang et al., 2023; Xu et al., 2024c; Huang et al., 2023) by providing factually incorrect information. Retrieval Augmented Generation (RAG) was proposed (Lewis et al., 2020; Guu et al., 2020) in order to address these issues as well as others and make LLMs more trustworthy.
The key idea behind RAG is to enhance the generative model’s capabilities by integrating a retrieval system that fetches relevant passages from a large corpus of data. In this setting, when a query is received, the retrieval system first identifies and retrieves pertinent information, which is fed into the generative model’s context for a more accurate and relevant response. Instead of the model storing information within its weights, RAG effectively leverages external knowledge, reducing hallucinations (by grounding the LLM reply in reliable sources), and ensuring that responses contain up-to-date knowledge (e.g., by accessing the Internet), all without requiring expensive training.
More specifically, there are two main stages in a RAG pipeline: data preparation and query execution. During data preparation, one constructs a vector database (DB) populated with embeddings and their corresponding data items such as documents. During query execution, one constructs an embedding of that query and retrieves data items in the store with similar embeddings.
Intense recent research efforts have been put into RAG (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2024; Huang & Huang, 2024; Yu et al., 2024; Mialon et al., 2023; Li et al., 2022). On one hand, different RAG designs have been proposed, for example RAPTOR (Sarthi et al., 2024), Self-RAG (Asai et al., 2023), Chain-of-Note (Yu et al., 2023), and many others (Abdallah & Jatowt, 2024; Delile et al., 2024; Edge et al., 2024; Manathunga & Illangasekara, 2023; Zeng et al., 2024; Wewer et al., 2021; Xu et al., 2024b). In general, these schemes focus on making the retrieved data more accurate and relevant to the query. There have also been efforts into benchmarking and datasets for RAG evaluation (Chen et al., 2024b; Xiong et al., 2024; Lyu et al., 2024; Es et al., 2023).
Despite all these advances, we observe that no existing RAG scheme or evaluation methodology explicitly targets an important class of problems that come with a high degree of multi-aspectuality. These are problems that require combining several (potentially many) significantly different aspects in a single query. As a simple illustrative example of such a query, consider the question “What car did Alexander the Great drive?”, and assume that the queried model has not been trained on history. When using RAG, to answer this question accurately, one would retrieve two documents, one describing Alexander the Great and one outlining the history of car manufacturing. However, the embeddings of these two documents could be far away from each other in the embedding space. At the same time, such queries are common in different industry settings, as indicated by extensive discussions with our industry collaborators. Imagine a chemical processing plant experiencing an equipment accident. One could use an LLM to find the accident cause, which might require the retrieval of multiple, potentially confidential documents to provide the necessary context. These documents could be related to different aspects, for example psychological profiles of workers (“Was the accident due to mismanaging a worker?”), equipment purchase records (“Was some equipment part too old?”), maintenance (“Was some equipment part rusty?”), weather (“Was there a particularly strong thunderstorm at the accident time that could have caused dangerous power spikes in the grid?”), or even microclimate (“Was it too humid for an extended period of time in the production hall?”). As we illustrate in Section 4, such problems pose challenges for existing RAG schemes and have been unaddressed by modern RAG benchmarking pipelines.
@@
In this work, we propose Multi-Head RAG (MRAG): a scheme that addresses the above problem. Common practice in modern RAG designs is the use of embeddings based on last-layer decoder block activations. Our key idea is to use instead the activations of the multi-head attention part of the decoder block as embeddings. The Transformer architecture can be seen as a pipeline with many (e.g., 96 for GPT-3 (Wang et al., 2024b)) blocks, where a single block consists of an attention module and a feed-forward module. Each individual attention module is multi-headed: it consists of multiple parts called heads that learn different sets of weight matrices; see Figure 1 for an overview. It is conjectured that these different heads could capture different aspects of the processed data. We use this as a driving design feature that facilitates capturing the potential multi-aspectuality of the data without increasing space requirements compared to standard RAG, and without any fine-tuning or other modifications to the harnessed model (contribution 1).
@@
Such multi-aspect embeddings are then directly used for both data items and query representation. Considering multi-aspectuality explicitly comes with challenges. For example, it is unclear how to assess whether a RAG solution does indeed harness multiple aspects when fetching documents. For this, we establish an evaluation methodology as well as a full data construction and query processing pipeline that implements the multi-aspect embedding idea (contribution 2). Our datasets facilitate broad evaluation by considering both fully-automatically generated, synthetic data as well as data based on specific industry use cases that show the benefits of MRAG (contribution 3). We ensure the relevance of our RAG datasets in real use cases by working directly with tech leaders (e.g., a generative AI division head) from 3 corporations, all of which actively use RAG in their own LLM infrastructures. Our evaluation illustrates the benefits in the relevance of retrieved documents, for example 20% over a modern RAG baseline for fetching multi-aspect Wikipedia articles, and comparable performance for single-aspect queries (contribution 4). We also show how MRAG and its benchmarking principles can be seamlessly integrated with both existing RAG solutions and benchmarking frameworks such as RAGAS (contribution 5). MRAG’s code is publicly available [^1] https://github.com/spcl/MRAG.
2 The MRAG Formulation & Pipeline
We now present in detail the mathematical underpinning of MRAG and its corresponding pipeline.
Decoder Formulation We first introduce formally the decoder architecture. We omit, for clarity, unnecessary details such as layer normalizations. The input is a text chunk that consists of $n$ tokens. The output of attention head $h$ for the $i$th token $x_{i}$ is defined as follows (Vaswani et al., 2017):
$\vspace{-0.5em}w_{ij}=\text{softmax}\left(\left(\mathbf{q}_{i}^{h}\right)^{T}\mathbf{k}_{j}^{h}\right),\quad\mathbf{q}_{i}^{h}=\mathbf{W}_{q}^{h}\mathbf{x}_{i},\quad\mathbf{k}_{j}^{h}=\mathbf{W}_{k}^{h}\mathbf{x}_{j},\quad\mathbf{v}_{j}^{h}=\mathbf{W}_{v}^{h}\mathbf{x}_{j}$ (2)
[…]
where matrix $\mathbf{W}_{o}$ is the linear layer that combines the outcomes of all the attention heads. This step is then followed by the Transformer feed-forward layer.
[…]
Multi-Head RAG Formulation The key idea behind MRAG is simple: instead of the single activation vector generated by the last feed-forward decoder layer for the last token, we harness the $H$ separate activation vectors generated by the last attention layer for the last token, before processing it via $\mathbf{W_{o}}$. This can be formulated as a set of embeddings $\mathcal{S}$:
$\mathcal{S}=\{\mathbf{e}_{k}\forall_{k}\}\quad\text{where}\quad\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})$ (4)
which is simply the set of all outputs from the attention heads on the last token $\mathbf{x}_{n}$ of the input. As processing with multiple heads does not change the size of the output vector, $\mathcal{S}$ has the same space requirements as standard RAG. However, because we capture the separate embeddings before their mixing with $\mathbf{W}_{o}$, we conjecture that it gives more information about what the different parts of the input attend to, facilitating capturing multi-aspectuality.
Naming We use the terms “single-aspect embedding” and “multi-aspect embedding” to refer to, respectively, a small embedding extracted from a single attention head and a collection of all single-aspect embeddings extracted from an attention layer.
2.1 Overview of the Multi-Head RAG Pipeline
We now describe how the above embedding model fits the RAG pipeline. Figure 2 shows a summary of the design. The MRAG pipeline consists of two main parts, dedicated to data preparation [Uncaptioned image] and query execution [Uncaptioned image]. Both parts heavily use the data store [Uncaptioned image] (vector DB).
2.1.1 Data Preparation
When preparing data [Uncaptioned image], we populate a data store [Uncaptioned image] with multi-aspect MRAG text embeddings [Uncaptioned image] and their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image] (MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data). We create the multi-aspect embedding [Uncaptioned image] of each text chunk [Uncaptioned image] using a selected decoder-based embedding model [Uncaptioned image] (this part is detailed in Section 2.2). The user of the pipeline can plug in their model [Uncaptioned image] of choice as well as use their input data. We also offer a dedicated synthetic data generator [Uncaptioned image] that can be used to construct multi-aspect input documents [Uncaptioned image] (we detail this part in Section 3) for evaluation purposes.
MRAG stores data differently than standard RAG, where a single embedding [Uncaptioned image] points to a single text chunk [Uncaptioned image]. For MRAG, each multi-aspect embedding consists of $h$ single-aspect embeddings [Uncaptioned image], each pointing to the original text chunk [Uncaptioned image]. So the data store [Uncaptioned image] contains $h$ embedding spaces, each capturing a different aspect of the text. This crucial feature allows MRAG to compare query [Uncaptioned image] and text chunks [Uncaptioned image] in multiple embedding spaces that capture multiple aspects of the data.
2.1.2 Query Execution
During query execution [Uncaptioned image], we first generate a multi-aspect embedding [Uncaptioned image] of the input query [Uncaptioned image], using the selected embedding model [Uncaptioned image] (details in Section 2.2). Then, we find the nearest multi-aspect embeddings [Uncaptioned image] and their corresponding text chunks [Uncaptioned image] in the data store [Uncaptioned image] using a special multi-aspect retrieval strategy [Uncaptioned image] (detailed in Section 2.3). We ensure that there is no overhead in latency due to multiple aspects because computing these different smaller embeddings is done fully in parallel. Finally, the retrieved data can optionally be assessed [Uncaptioned image] with novel metrics regarding how well it corresponds to the multi-aspect requirements (detailed in Section 3). As with the data preparation [Uncaptioned image] stage, the query execution [Uncaptioned image] stage is flexible, and the user can plug in their models [Uncaptioned image] / [Uncaptioned image] of choice and use their own queries [Uncaptioned image]. We also offer a dedicated synthetic query generator [Uncaptioned image] that can be used to construct multi-aspect input queries [Uncaptioned image] (detailed in Section 3) for evaluation purposes.

### v5

@@ Abstract
Retrieval Augmented Generation (RAG) enhances the abilities of Large Language Models (LLMs) by enabling the retrieval of documents into the LLM context to provide more accurate and relevant responses. Existing RAG solutions do not focus on queries that may require fetching multiple documents with substantially different contents. Such queries occur frequently, but are challenging because the embeddings of these documents may be distant in the embedding space, making it hard to retrieve them all. This paper introduces Multi-Head RAG (MRAG), a novel scheme designed to address this gap with a simple yet powerful idea: leveraging activations of Transformer’s multi-head attention layer, instead of the decoder layer, as keys for fetching multi-aspect documents. The driving observation is that different attention heads learn to capture different data aspects. Harnessing the corresponding activations results in embeddings that represent various facets of data items and queries, improving the retrieval accuracy for complex queries. We provide an evaluation methodology and metrics, multi-aspect datasets, and real-world use cases to demonstrate MRAG’s effectiveness. We show MRAG’s design advantages over 18 RAG baselines, empirical improvements of up to 20% in retrieval success ratios, and benefits for downstream LLM generation. MRAG can be seamlessly integrated with existing RAG frameworks and benchmarks.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020) emerged as a promising remedy for several key limitations of Large Language Models (LLMs). By decoupling knowledge from model weights, RAG reduces the risk of leaking confidential data (Yan et al., 2025; Shi et al., 2024; Patil et al., 2024), a critical concern when training on sensitive corpora. It also mitigates hallucinations (Zhang et al., 2025; Xu et al., 2025; Huang et al., 2025b) by grounding LLM outputs in retrieved, verifiable information. The core mechanism involves augmenting a generative LLM with a retrieval module that fetches relevant passages from an external corpus in response to a query. Rather than relying solely on static, parametric knowledge, RAG dynamically incorporates retrieved content into the model’s context, enabling more accurate and up-to-date responses. While early RAG systems required training task-specific retrievers and readers (Humeau et al., 2020; Singh et al., 2021), the current trend favors lightweight, in-context learning (ICL) approaches (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2025; Huang & Huang, 2024), which avoid the cost and complexity of retraining and allow for rapid knowledge updates without modifying the underlying LLM parameters.
A RAG pipeline consists of two main stages: data preparation and query execution. In the preparation stage, a vector database (DB) is constructed by embedding a collection of documents and storing these embeddings alongside their associated content. At inference time, the query is similarly embedded, and nearest-neighbor search retrieves the most relevant data items, which are then passed to the LLM for final answer generation. Ongoing developments in RAG have led to variety of different RAG designs (Gao et al., 2024; Zhao et al., 2024; Hu & Lu, 2025; Huang & Huang, 2024; Yu et al., 2024a; Mialon et al., 2023; Li et al., 2022; Chen et al., 2024c; Xiong et al., 2024; Lyu et al., 2025; Es et al., 2024; Sarthi et al., 2024; Asai et al., 2024; Abdallah & Jatowt, 2024; Delile et al., 2024; Edge et al., 2025; Manathunga & Illangasekara, 2023; Zeng et al., 2024; Wewer et al., 2021; Xu et al., 2024; Huang et al., 2025a; Yang et al., 2025; Su et al., 2025; Zhao et al., 2025a; b; Chen et al., 2024d; Chan et al., 2024; Jiang et al., 2024; Yu et al., 2024b).
Yet, no existing RAG method or benchmark explicitly targets multi-aspectual problems, that is, queries requiring the integration of multiple, semantically distinct aspects. For example, answering “What car did Alexander the Great drive?” (assuming no historical pretraining) requires retrieving unrelated documents on Alexander the Great and on car manufacturing, whose embeddings may lie far apart in the vector space. Such multi-aspect queries are common in industrial settings, as confirmed by extensive discussions with our industry collaborators and further supported by our analysis of over 35 industry reports (Office of Railroad Safety, 2014; Columbia Accident Investigation Board, 2003; Zografos et al., 2013; Reason et al., 2006; Kalia, 2013; Packham, 2017; Messiou et al., 2023; Multer et al., 2013; Coury et al., 2010; Harle, 1997; O’Hare, 2000; Gordon et al., 2005; Bridger, 2021; OpenText, 2022) (details are in Appendix A; we considered accident prevention, healthcare, airport management, and others). For example, in a chemical plant accident, determining the cause might require accessing diverse and confidential documents related to worker psychology (“Was it mismanagement?”), equipment records (“Was a part outdated or rusty?”), weather conditions (“Were there power spikes due to a storm?”), or even microclimate (“Was prolonged humidity a factor?”). As shown in Section 5, such cases have been unaddressed by modern RAG schemes and benchmarks (Chen et al., 2024c; Xiong et al., 2024; Lyu et al., 2025; Es et al., 2024).
@@
In this work, we propose Multi-Head RAG (MRAG): a scheme that addresses the above problem (contribution 1). Common practice in many modern RAG designs is to use embeddings derived from last-layer decoder block activations of a decoder-based embedding LLM (a language model specifically fine-tuned to provide high-quality embeddings). Examples of such embedding models are SFR-Embedding-Mistral and E5-Mistral-7B. Our key idea is to extend this design by incorporating activations from the multi-head attention (MHA) modules of decoder blocks as embedding sources. This enables the representation of multiple distinct aspects of the input text. Specifically, a Transformer consists of a stack of blocks (e.g., 96 in GPT-3 (Wang et al., 2025b)), each containing an MHA module with multiple heads that are trained with separate parameter sets. Through a survey of literature into Transformer design and interpretability, we find empirical and theoretical support for the conjecture that different heads specialize in different aspects of the input (details are in Section 2.4). This enables efficient multi-aspect embeddings without increasing space or compute costs compared to standard RAG, and without requiring any additional fine-tuning or architectural modifications to the base model.
@@
Considering multi-aspectuality comes with challenges. For example, it is unclear how to assess whether a RAG solution does indeed harness multiple aspects when fetching documents. For this, we develop a multi-aspect RAG pipeline that includes data preparation and query processing with both multi-aspect retrieval and ranking schemes (contribution 2). We also establish an evaluation methodology and provide multi-aspect datasets, complementing existing RAG benchmarks (Chen et al., 2024c; Xiong et al., 2024; Lyu et al., 2025; Es et al., 2024) (contribution 3). We ensure the relevance of our RAG datasets in real use cases by working directly with tech leaders (e.g., a generative AI division head) from 3 corporations, all of which actively use RAG in their own LLM infrastructures. We illustrate the advantages of MRAG over 18 traditional and modern RAG designs in various design criteria and in both time and space complexities (contribution 4). In evaluation, MRAG enhances the relevance of retrieved documents by up to 20% over modern RAG baselines, offers comparable performance without degradation for single-aspect queries, and benefits the downstream LLM generation (contribution 5). Thanks to its simplicity, MRAG can be seamlessly integrated into any stores while its benchmarking methodology can straightforwardly extend benchmarks such as RAGAs (contribution 6). MRAG’s code is publicly available [^1] https://github.com/spcl/MRAG.
2 MRAG: Design & Implementation
A typical RAG scheme (see Figure 2) consists of two main parts: data preprocessing [Uncaptioned image] and query execution [Uncaptioned image]; both parts heavily use an embedding model [Uncaptioned image] and the data store [Uncaptioned image]. During preprocessing, each document in the database is encoded into one or more embeddings using the embedding model; these embeddings are stored with that document (usually as a key-value pair in a vector DB). During query execution, the embedding of the user-provided query is constructed using the same embedding model; then, the retriever fetches candidate documents based on embedding similarity, the reranker refines their ordering using more precise scoring, and the reader uses a downstream generative model to synthesize the final answer from the top-ranked documents.
@@ 2.1 Constructing Multi-Aspect Embeddings [Uncaptioned image]
An embedding is constructed for each data item in a pre-existing database (part [Uncaptioned image]) and for the user query (part [Uncaptioned image]). In standard modern RAG, given an input chunk of $n$ tokens constituting a document or chunk to be embedded, one obtains a corresponding embedding by applying the embedding model to the data item and extracting the activation of the final feed-forward layer for the last token $\mathbf{x}_{n}$, i.e., $\texttt{feed-forward}(\texttt{multi-head}(\mathbf{x}_{n}))\in\mathbb{R}^{d}$. Appendix B provides additional mathematical details.
In MRAG, instead of relying on this single embedding, the activations of all $h$ attention heads are obtained before they are merged by the final projection layer. Specifically, for the last token $\mathbf{x}_{n}$, we extract the set of head-specific vectors $\mathcal{S}=\{\mathbf{e}_{k}\}_{k=1}^{h}$, where each $\mathbf{e}_{k}=\text{head}^{k}(\mathbf{x}_{n})\in\mathbb{R}^{d/h}$ is referred to as a “single-aspect embedding” and $\mathcal{S}$ is called a “multi-aspect embedding”. This provides $h$ semantically diverse embeddings per input, reflecting the different perspectives captured by each attention head. Since these vectors are extracted from the same internal activation used by standard RAG, the overall space and time requirements remain unchanged. However, the explicit separation of heads allows MRAG to preserve fine-grained aspectual distinctions, facilitating multi-aspect representation.
2.2 Data Preprocessing [Uncaptioned image]
We populate a data store [Uncaptioned image] with multi-aspect embeddings [Uncaptioned image] for their corresponding documents [Uncaptioned image] or text chunks [Uncaptioned image]. MRAG is orthogonal to the type of data being embedded, and while we primarily use chunking of documents in order to reflect modern RAG pipelines, one can also embed whole documents or even other types of data. Unlike in standard RAG, where a single embedding [Uncaptioned image] points to a single text chunk [Uncaptioned image], in MRAG, each out of $h$ single-aspect embeddings [Uncaptioned image] points to the original text chunk [Uncaptioned image] (i.e., the data store [Uncaptioned image] contains $h$ embedding spaces). This crucial feature allows MRAG to compare query [Uncaptioned image] and text chunks [Uncaptioned image] in multiple embedding spaces that capture multiple aspects of the data.
MRAG performs one-time importance scoring for each embedding space $i$, capturing its relevance to the dataset. Each score $s_{i}$ combines: (i) the average L2 norm $a_{i}$ of embeddings in space $i$, reflecting attention strength, and (ii) the average cosine distance $b_{i}$ between sampled embeddings, approximating its semantic spread. The final score is $s_{i}=a_{i}\cdot b_{i}$, encouraging both head relevance (through high $a_{i}$) and high representational diversity (through high $b_{i}$). Full details are provided in Algorithm 1, Appendix C.1.
2.3 Query Execution
During query execution, MRAG first generates a multi-aspect embedding $\mathcal{S}$ of the input query (cf. Section 2.1). These embeddings are computed fully in parallel during the same inference pass, so this multi-vector representation introduces no additional latency.
Retriever. Given $\mathcal{S}$, the retriever conducts parallel retrieval of the top-$c$ nearest chunks within each embedding space, and the aggregation of the resulting candidates across all $h$ spaces.
Reranker. Once $hc$ candidate chunks are retrieved across all heads, the ranker stage consolidates them into a single list of top-$k$ results using a simple but effective voting strategy. For each candidate chunk at position $p$ in the ranked list for space $i$, we assign a score of $s_{i}\cdot 2^{-p}$, where $s_{i}$ is the precomputed importance score. This exponentially discounts lower-ranked candidates and balances influence across heads. The final top-$k$ list is obtained by globally sorting all candidates by these scores. The voting procedure is described in Algorithm 2 in Appendix C.2.
Reader. The top-$k$ retrieved results are inserted into the LLM context using a multi-aspect prompt template (prompts are fully specified in Appendix D). Each result is placed in a separate section of the prompt. Stored metadata can be included alongside each chunk to provide additional context, such as the aspect or the chunk category.
2.4 Capturing Multi-Aspectuality Without Additional Training
A key design decision in MRAG is to use the hidden representations immediately after the attention block in the last decoder layer, without additional fine-tuning, to avoid training overhead and make deployment easy. This is motivated by growing evidence that attention heads in Transformer models naturally converge during training to focus on distinct aspects of the input. For example, Wang et al. (2025a) show that heads diverge into clusters specialized for different input patterns, while Olsson et al. (2022) and others (McDougall et al., 2024; Gould et al., 2024) discover that heads focus on repeated sequences or named entities. Similar trends are observed in BERT (Clark et al., 2019; Kovaleva et al., 2019; Htut et al., 2019). We provide more details in a brief literature survey (Appendix E.1). Given these findings, we assume that even without additional fine-tuning, the embeddings output by MHA already encode multi-aspectuality suitable for downstream retrieval. Our own brief attention pattern analysis (Appendix E.2) confirm shifting token focus across heads.
2.5 Parallelization and Systems Considerations
MRAG evenly distributes the embedding dimensionality across attention heads, keeping total storage and compute $O(nd)$ (same as standard RAG; we show a more detailed complexity analysis in the following section). It uses off-the-shelf ANN indexes that support parallel subspace search, and it avoids dynamic or variable-length vectors that complicate indexing and caching.
Modern vector databases already support the parallel multi-vector search required by our approach. Milvus can execute distributed, parallel ANN queries across shards with low latency (Wang et al., 2025c; Clavié et al., 2024; Wang et al., 2021). Pinecone provides cascading retrieval with fixed-size multi-vector encodings (such as ConstBERT) to maintain predictable costs. Furthermore, ESPN tackles multi-vector retrieval at SSD and GPU levels by implementing prefetching pipelines and storage bypass mechanisms, achieving near-memory query latency, even when dealing with large indices (Shrestha et al., 2024). Taken together, these technologies illustrate that multi-vector retrieval, especially when using fixed-size vectors as MRAG does, can be implemented at scale with negligible practical overhead.
3 Compute & Storage Complexity Analysis
We analyze the runtime and space complexity of MRAG and 18 RAG baselines, the results are in Table 1 (the runtime metrics of work/depth and the notation are explained in the table caption). Derivation details and a broader discussion are in Appendix F. Overall, MRAG achieves competitive results in all aspects. At inference time, it extracts $h$ attention-head embeddings in parallel from a single forward pass, resulting in the same latency as standard RAG. Preprocessing is lightweight, requiring only one pass per document and simple statistics for head scoring, avoiding the cost of training or complex structure construction. Storage overhead is minimal, as $h$ single-aspect embeddings per data item have the same dimensionality as a standard embedding. In contrast, prior schemes like Poly-encoder (Humeau et al., 2020) and ColBERT (Khattab & Zaharia, 2020) incur significant cost due to using many token-level embeddings or time-consuming training rounds.
Table 1: Time and storage complexity of different RAG schemes; MRAG is the only scheme to match the results of a plain vanilla RAG. Work is the total number of all operations, Depth is runtime complexity assuming unlimited parallel processing units; both are established measures of analyzing parallel algorithms. Retrieval describes steps during inference, including any postprocessing before returning documents or summaries. Preprocessing includes all steps before inference. $n$: number of data items in a database (i.e., document chunks), $d$: embedding dimensionality, $k$: number of retrieved top chunks per single user query, $l_{q}$: average token length of the query, $l_{d}$: average token length of the document, $W_{m}$$D_{m}$: work/depth to run a transformer-based model, $W_{e}$$D_{e}$: work/depth to embed a graph, $W_{i}$$D_{i}$: work/depth to index a graph, $s$: polynomial function that models the complexity of various additional scheme-specific design decisions, heuristics, etc., which cannot be straightforwardly modeled with closed-form expressions (e.g., count of graph communities, toy-graph size, self-note length, BM25 matching cost, keyword matching cost); it scales at most linearly with $n$ and is typically considerably larger than $O(1)$.
Scheme Retrieval Preprocessing Storage
Work Depth Work Depth
Vanilla RAG $O(W_{m}+nd)$ $O(D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
Poly-encoders (Humeau et al., 2020) $O(W_{m}+nd+s(n)d)$ $O(D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
Lewis et al. (2020) $O(2W_{m}+nd)$ $O(2D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
ColBERT (Khattab & Zaharia, 2020) $O(W_{m}+ndl_{q}l_{d})$ $O(D_{m}+\log dl_{d})$ $O(W_{m}n)$ $O(D_{m})$ $O(ndl_{d})$
EMDR (Singh et al., 2021) $O((k+1)W_{m}+nd)$ $O(2D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
Self-RAG (Asai et al., 2024) $O((2k+1)W_{m}+nd+k)$ $O(2D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
Chain-of-Note (Yu et al., 2024b) $O((ks(n)+1)W_{m}+nd)$ $O((ks(n)+1)D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
RAPTOR (Sarthi et al., 2024) $O(W_{m}+nd+s(n)d)$ $O(D_{m}+\log d)$ $O(W_{m}n+W_{m}s(n))$ $O(D_{m})$ $O(nd+s(n)d)$
RAGraph (Jiang et al., 2024) $O(W_{m}+nd+s(n)d)$ $O(D_{m}+\log d)$ $O(W_{m}+W_{e})$ $O(D_{m}+D_{e})$ $O(nd+s(n)d)$
RQ-RAG (Chan et al., 2024) $O(s(n)(W_{m}+nd))$ $O(s(n)(D_{m}+\log d))$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
ActiveRAG (Xu et al., 2024) $O(3W_{m}+nd)$ $O(2D_{m}+\log n)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
HiQA (Chen et al., 2024d) $O(W_{m}+nd+s(n)d)$ $O(D_{m}+\log d)$ $O(W_{m}n+s(n)d)$ $O(D_{m})$ $O(nd+s(n)d)$
GraphRAG (Edge et al., 2025) $O(s(n)(W_{m}+d))$ $O(D_{m}+s(n)d)$ $O(W_{m}n+W_{m}s(n)+W_{e})$ $O(D_{m}+D_{e})$ $O(nd+s(n)d)$
Fusion RAG (Rackauckas, 2024) $O(s(n)W_{m}+knd)$ $O(2D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
Meta-chunking (Zhao et al., 2025b) $O(W_{m}+(n+s(n))d)$ $O(D_{m}+\log d)$ $O(l_{d}W_{m}+s(n)W_{m})$ $O((l_{d}+1)D_{m})$ $O(nd+s(n)d)$
MoC (Zhao et al., 2025a) $O(W_{m}+(n+s(n))d)$ $O(D_{m}+\log d)$ $O(W_{m}n+s(n))$ $O(D_{m})$ $O(nd+s(n)d)$
Parametric RAG (Su et al., 2025) $O(W_{m}+nd+s(n))$ $O(D_{m}+\log d+s(n))$ $O(W_{m}ns(n))$ $O(D_{m}s(n))$ $O(ns(n))$
SuperRAG (Yang et al., 2025) $O(W_{m}+nd+W_{i})$ $O(D_{m}+\log d+D_{i})$ $O(W_{m}n+s(n)+W_{e})$ $O(D_{m}+D_{e})$ $O(nd+s(n))$
HiRAG (Huang et al., 2025a) $O(W_{m}+nd+W_{i})$ $O(D_{m}+\log d+D_{i})$ $O(W_{m}n+s(n)+W_{e})$ $O(D_{m}+D_{e})$ $O(nd+s(n))$
MRAG [This Work] $O(W_{m}+nd)$ $O(D_{m}+\log d)$ $O(W_{m}n)$ $O(D_{m})$ $O(nd)$
4 Benchmarking Multi-Aspectuality
To assess how well MRAG performs on multi-aspect queries, we need (1) datasets of multi-aspectual documents, (2) queries to the LLM that require retrieving multi-aspect documents, and (3) metrics that assess how well a RAG scheme retrieves such multi-aspect data. We now summarize these three elements; details are in Appendix G.
We construct three multi-aspect datasets: (1) a synthetic Wikipedia dataset with documents sampled from clearly distinct categories (e.g., countries, shipwrecks, board games); (2) a real-world-based legal document dataset annotated by a legal areas (energy law, family law, criminal law, etc.) or document language style (aggressive, mild, neutral, etc.); and (3) a real-world-based dataset of industry accident reports, categorized by cause. For each dataset, we generate multi-aspect queries using GPT-4o, combining $n$ distinct categories into single queries (with $n\in\{1,2,3,4,5,6,10,15,20,25\}$). For example, a query with 10 aspects must contain a question about 10 different documents from 10 different categories.
To evaluate retrieval performance, we introduce three metrics for assessing multi-aspectuality. Let $Q$ be a query, $S$ a Reranker scheme, and $Q_{rel}$ the ideal set of $n$ relevant documents. The Retrieval Success Ratio is defined as $\Xi(Q,n)=\frac{|S(Q,n)\cap Q_{rel}|}{|Q_{rel}|}$, measuring the fraction of exactly matched documents. The Category Retrieval Success Ratio $\Xi_{c}$ extends this by also accepting matches from the same categories as those in $Q_{rel}$ even if the exact document has not been matched. Finally, we define a tunable Weighted Success Ratio $\Xi_{w}=\frac{w\cdot\Xi+\Xi_{c}}{w+1}$, allowing the user to adjust the trade-off between exact and category-level matches via the weight $w$.
We motivate category-level retrieval using both previous works (V et al., 2025; Liu et al., 2021; Chen et al., 2024a; Xiao et al., 2024) as well as our hands-on experience in legal and industrial accident analysis. Namely, even when exact matches are missing, retrieving documents from the same semantic category can enhance generation, as supported by the classic clustering hypothesis in information retrieval (V et al., 2025; Liu et al., 2021): documents in the same “neighborhood” are likely to be relevant to the same query. Recent studies further validate this across open-domain QA and ontology-guided RAG (Chen et al., 2024a; Xiao et al., 2024).
5 Evaluation
We now illustrate the advantages of MRAG over the state of the art. Further details on evaluation setup and additional results are in – respectively – Appendix H and I.
Comparison Baselines. We consider three main baselines: Standard RAG, Split RAG, and Fusion RAG (Rackauckas, 2024). The first represents a modern RAG pipeline in which each document uses the activations of the last decoder layer as its embedding. The second is a blend between Standard RAG and MRAG; it splits the activation of the last decoder layer in the same way as MRAG and applies a voting strategy. The purpose of Split RAG is to show that MRAG’s benefits come from using the multi-head output as embedding and not merely using multiple embedding spaces. Additionally, we consider Fusion RAG (Rackauckas, 2024), an optional mechanism that we harness to further enhance the benefits of MRAG at the tradeoff of additional tokens (detailed in Section 5.3).